In triangle ABC, D is the midpoint of line AB and E is the midpoint of line BC. If AC= 3x-15 and DE= 6, what is the value of x?
Answer by edjones(8007)

The triangles ABC and DEB are similar. The sides of ABC are twice as long as DEB.
3x-15=2*6
3x-15=12
3x=27
x=9
.
Ed